OpenStudy (anonymous):

expresss complex number in rectangular form \[\sqrt{7}*cis*2.1 \]

OpenStudy (anonymous):

is that cosine ?

OpenStudy (anonymous):

its cis cos+isin

OpenStudy (anonymous):

yes it is stupid notation for \[cos(\theta)+i sin(\theta)\]

OpenStudy (anonymous):

im trying square of 7 which is 2.64 times cos but cos(?)

OpenStudy (anonymous):

i also assume that you are working in radians, not degrees. So just use a calculator. \[\sqrt{7} [cos(2.1)+i sin(2.1)]\]

OpenStudy (anonymous):

i get -1.336 for a and 2.284 for b (rounded) so answer is -1.336 + 2.284 i

OpenStudy (amistre64):

the cosine part relates to the x coord; the i sin part relates to the y coord right?

OpenStudy (amistre64):

Did the others parse it tight to be: cos(2.1) + i sin(2.1)?

OpenStudy (amistre64):

(-1.34, 42.30) is what i get with radian measurements...

OpenStudy (amistre64):

2.28.. musta hit a wrong buttononthe calc :)

OpenStudy (anonymous):

yes cosine is x and sine is y, usually written as a and b as in a + bi. I doubt it is (-1.34,42.3) for two reasons: the rectangular form of a complex number is a number, not a coordinate. so it should look like a + bi also we know that the absolute value of that number is \[\sqrt7\]

OpenStudy (amistre64):

correct, its similar to polar coordinates in that respect. But i think the exercise here is to see that complex numbers are nothing new and that they can be plotted in the same manner as rectangular coords...right?